OpenStudy (zoiedaniella):

Can someone help me understand this and answer it? What is the simplified form of the quantity of x plus 6, all over the quantity of x plus 4 + the quantity of x minus 3, all over 3?

OpenStudy (zoiedaniella):

Which is: X+6 X-3 _____ ------ X+4 + 3

OpenStudy (mjdennis):

This is like adding (1/3) + (1/4), you have to convert the fractions before adding them together:

OpenStudy (zoiedaniella):

Do you mean multiplying the denominators so they're the same?

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (mjdennis):

My network is dropping my posts. Yes, so take each fraction. 1/3, for instance. Multiply top and bottom by the other denominator. (1/3) * (4/4) = 4 /12 The other term in my example becomes 3/12, and now you can add 4/12 + 3/12

OpenStudy (mjdennis):

The first term is easy, right? X+6 (3) _____ ___ X+4 (3)

OpenStudy (mjdennis):

What do you get for the second term? Remember to multiply top and bottom by the bottom of the other term.

OpenStudy (zoiedaniella):

Sorry my internet is being a pain too. So the first term would be 3x+18 ------ 3x+12 and then I believe the 2nd term would be x-3 (x+4)= x^2+x-12 ---------- 3x+12 If thats correct.

OpenStudy (zoiedaniella):

is it right to have a trinomial as the numerator?

OpenStudy (zoiedaniella):

3x+18 x^2+x-12 ------ + ---------- 3x+12 3x+12 and then I would just combine the top 2 right? How would I combine that tho

OpenStudy (mjdennis):

Yes it is!

OpenStudy (phi):

now add the tops

OpenStudy (phi):

btw, I would leave the bottom as 3(x+4) it's easier to see if anything will "cancel" if you do that

OpenStudy (mjdennis):

well, you combine the fractions just like adding (4/12) +(3/12) = (4+3)/12 = 7/12 Or, are you having trouble adding the two top expressions? To do this, we 'group like terms together', a math code for group any terms that have x^2 together, then x terms, then numbers. So if you had 4x+16 + 3x^2+2x-1 you have 3x^2 +4x+2x +16-1 And they add to 3x^2 +6x +15

OpenStudy (zoiedaniella):

I think I got it, it would be x^2 +4x+6 ---------- 3(X+4) Thank you so much!! I just totally forgot how to do it lol. Will medal and fan.

OpenStudy (mjdennis):

Last step: factor the top, if you can.

OpenStudy (mjdennis):

Oh, and pretend I used the correct math terms, "numerator" and "denominator", instead of "top" and "bottom"